A car is travelling at a constant speed of 11m/s. Passes a trooper hidden behind a Billboard one second of the car passes the Billboard, the trooper sets out from the Billboard to catch it, acelerating at a constant speed of 22m/s².how long does it take her to take the car?
1
Expert's answer
2020-02-21T10:03:40-0500

Let the time of motion of the trooper is t, then distance traveled by her


"d=0.5at^2"

The distance traveled by a car


"d=v(t+1)"

"v(t+1)=0.5at^2\to 11(t+1)=0.5(22)t^2"

"t=1.6\ s"